OpenStudy (anonymous):

Tara's car accelerated from 10 miles per hour to 60 miles per hour in 12 seconds. What is the approximate rate of change in the miles per hour that Tara's car accelerated? A) 4 miles per hour per second B) 5 miles per hour per second C) 18 miles per hour per second D) 48 miles per hour per second

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Amber_kinG

OpenStudy (anonymous):



OpenStudy (anonymous):

this site sucks everyone goto brainly.com FOR INSTENT ANSWERS!

OpenStudy (michele_laino):

the requested acceleration, is: \[\large a = \frac{{{v_{final}} - {v_{initial}}}}{{\Delta t}} = \frac{{60 - 10}}{{12}} = ...?{\text{miles/hours/seconds}}\]